A Bernoulli Trials experiment has p=15/23 probability of success on each trial.

What is the expected number of successes in five trials?

What is the expected number of successes in 16 trials?

What is the expected number of failures in 69 trials?

Enter your answers as whole numbers or FRACTIONS in lowest terms.
Answer #1

Solution :

Given that,

p = 15 / 23 = 0.65

q = 1 - p = 1 - 0.65 = 0.35

1) n = 5

expected value = = n * p = 5 * 0.65 = 3.25

2) n = 16

expected value = = n * p = 16 * 0.65 = 10.4

3) n = 69

expected value = = n * q = 69 * 0.35 = 24.15